# Articulation Work and Tinkering for Fairness in Machine Learning

### v1

@@ 5.3. Openness to the SOI paradigm and tinkering
We conclude our findings by pointing out some interview participants’ micro-strategies and attempts to integrate some SOI perspectives in their everyday research practices.
@@
Some of our interview participants, notably the two professors, R4 and R5, both situated in German universities along with R8, the professor from an Italian university, and R10, a professor from a UK university, expressed their wish for fair AI research to become more inclusive of SOI perspectives. Some of them acknowledged their own agency and responsibility in attaining this. In particular, R10 emphasized how he tried to negotiate and shape the conditions of his CS lab to align with a more inclusive research approach. This, as he explained, included actively taking responsibility for conducting more research with interdisciplinary actors, such as scholars from the social sciences. As a senior researcher, he had experienced that such collaborations did not emerge spontaneously or by themselves but required proactive efforts by researchers, such as to “negotiate and have a more active role in your occupation to be able to keep the things that you’re happy with.” (R10)
@@
Some interview participants built their own accounts of doing SOI work by gradually deviating from usually performed CS tasks. Drawing on Fujimura (1987) and Knorr (1979), we understand these micro-strategies as tinkering. We especially observed tinkering with SOI methods. For instance, R7, the professor at the Dutch University engaged his PhD students in conducting “short ethnographic studies”, as he coined them, at field sites of their collaborative partners. At the same time, the ethnographic study was further adapted and tinkered with, so that articulation work did not increase significantly. The professor conveyed that in his lab, ethnography - which is usually conducted for a longer period of time, with the researcher becoming very familiar with the field, following rigorous strategies of self-reflection, ‘thick description’, and a back-and-forth between theory and empirical material (Hammersley and Atkinson, [n. d.]; Spradley, [n. d.]; Geertz, [n. d.]) - was only conducted for several days, and without formal training. In another example, R1 told us a story about how he aimed to conduct a qualitative study to test users’ perceptions of an algorithmic system he and his colleagues were designing. However, to avoid additional articulation work (e.g. setting up a sampling strategy, reaching out to interview participants, organizing the interview setting, etc.), he turned to his colleagues as proxies for the intended users of the algorithmic system.
@@
Our findings show that there was a desire to accomplish impactful (fairness) research that coexisted with varying degrees of integration of SOI into everyday research practices. Even researchers with a more critical stance towards SOI demonstrated some (surface-level) engagement with it. Yet, due to the organizational constraints of the CS lab, SOI was adjusted or modified, e.g. tinkered with, to sidestep additional articulation work.
@@ 6. Discussion
Bringing Fujimura’s study in dialogue with our findings, we now discuss the gap between CS and SOI as a problem of ‘organizational alignment’.
@@
Our first argument is that computational researchers focus on the CS paradigm of fairness largely “because it works”, as to quote Knorr-Cetina ([n. d.]). On the contrary, adopting the SOI paradigm would necessitate acquiring new material resources and staff, reconfiguring publication venues, and dedicating additional time to novel methodological processes.
@@
In our study, we identified similar resources as crucial for making the CS paradigm of fairness research work. Much like in Fujimura’s observations, computational researchers pragmatically leveraged comparable resources for fairness research just as they would for other research problems. This included acquiring funding and utilizing technologies, staff and skills similar to how they approached “other ML problems” (R5). What partly set fairness research apart from other ML problems was its particular reliance on societal data that presented compelling societal cases (or ‘real-world scenarios’). Computational researchers encountered challenges in legitimizing fairness research without access to such data, highlighting that technical resources shaped what was (still) considered a meaningful advancement in the fair AI field.
@@
As Fujimura (1987) highlights, uncertainty increases articulation work as it constrains researchers’ possibilities to plan. In our case study, uncertainties posed constraints on computational researchers. The ‘projectification’ of research, with limited time frames, created general challenges for researchers in planning for the long-term. Information overload (e.g. ‘noise’) added to these uncertainties, as researchers were challenged to decide on their own accord which contributions to the field were important enough to extend, incorporate or reference in their work. In that regard, the lack of regulatory and technical standards (e.g. benchmarks, toolkits, agreed upon metrics) for fairness also increased uncertainty. As a result, it required more articulation work to find individualized solutions with the collaborative partners of computational researchers.
@@
If uncertainties and burdensome articulation work prevented computational researchers from fully engaging with SOI, how can SOI become a ‘doable problem’ for them? From our theoretical standpoint, we now examine this question for each of the three organizational levels.
@@
The doability of a research problem greatly relies on the positions of the social worlds. Fujimura (1987) and Clarke and Star (2008) have noted that at times, there may be multiple intersecting social worlds, each with its own interests. This can increase the complexity of the articulation work required. As previously discussed, the fair AI world faces new emerging actors, such as regulatory bodies, becoming more important. Furthermore, different interpretations within social worlds existed regarding what is considered doable. This does not necessarily imply that fairness loses credibility because it is “empty” (Laclau, [n. d.]; MacKillop, [n. d.]; Kalluri, [n. d.]), but that it can rather hold varying significance for different individuals (e.g., a so-called ‘boundary object’ (Star and Griesemer, [n. d.])). From this perspective, the much-requested call for standards, as sometimes indicated by our respondents, could potentially hinder interdisciplinary collaborations and alignment with diverse social worlds. Perhaps instead of such ethical or regulatory standards on fairness for computational researchers to translate into local (lab) practices, we need more alignment between different social worlds (e.g. CS community, regulatory bodies, companies, marginalized communities). This could include explicit stances from the social worlds’ actors acknowledging the SOI paradigm of fairness and making the SOI perspective an integral part of what is deemed as “success” (Knorr-Cetina, [n. d.]).
@@
On the level of the laboratory, the criteria of how research is evaluated, what counts as a successful publication, and which venues are treated as important, could be re-evaluated. The success of research is still mostly defined via publications (and within these, the ‘weight’ relies on rigid and decontextualized hierarchies of venue quality). The quest for other forms of visibility is, and will remain, an open problem for both scientometrics theory and academic institutions’ practices. Academic institutions’ ongoing developments of evaluation and indicator systems (which are, to some extent, inherently self-defeating) continues141414e.g.,KU Leuven’s position on the responsible use of metrics in the assessment of research(ers).. One potential approach to follow is the inclusion of meaningful SOI related goals into project proposals, so that project timelines explicitly account for research initiatives such as thorough user studies. To the same effect, it is also important to account for the fact that an enhanced SOI focus of a project will likely also require an enhanced role for SOI actors within the project.
@@
We have made the point that much of what drives the current situation is a need to limit articulation work for overburdened (computational) researchers. Therefore, we acknowledge that suggesting changes on the individual or experimental level could actually increase articulation work. However, we need alternative strategies that individual researchers can take if they wish to impact the kind of research that gets done. One strategy that we propose is taking a conscious awareness of how we allocate time to various tasks and a thoughtful consideration of how tasks are driven by optimizing productivity and efficiency. From our own experience in collaborating for this paper, we would like to emphasize that interdisciplinary alignment takes a considerable amount of time and that there are misses along the way. To navigate such phases successfully, requires, as one of our interview partners also pointed out, to allow our research to be guided by excitement and curiosity even if that initially seems undoable.
@@
Paradoxically, making SOI doable could also accentuate, or create new, hierarchies between paradigms, and disciplines. Resources are inherently unequal across disciplines and academic institutions (Merton, [n. d.]; Larregue, [n. d.]; Albert et al., [n. d.]). As supported by our findings, this can lead to a scenario where SOI methods and theories are gradually absorbed by computational disciplines that are already situated in a more powerful and financially advantaged position, particularly when it comes to ML research projects. In configuring interdisciplinary projects, close attention to fair resource and influence distributions amongst disciplines and institutions is required. Nevertheless, there is a need to address open questions about how to make something doable without perpetuating these hierarchies.
@@ 7. Limitations and Future Research
We hope that our study opens some novel avenues for investigating the evolvement of fair AI from the perspective of the CS researchers involved in it. Yet, our study is limited to only a small group of researchers. While we were especially interested in the perspectives and experiences of these computational researchers, this can reduce the scope of our work, and hence limit the generalizability of the claims made regarding CS laboratories at large. Future research could integrate the perspectives of other relevant actors in the laboratory. Beyond CS, the doability of fair AI depends on the collaborative efforts of engineers, data scientists, ethicists, research group leaders or IT experts who may be all situated in the (same) laboratory. Further research could investigate these micro-level actors’ interactions with each other and meso-level actors more closely.
@@
Second, we would find it interesting to further systematically explore our interviews regarding our respondents’ definitions of “fair AI” compared to a rather more flexible and open definition adopted while talking about their ongoing research projects and professional routines. Related to that, we were explicitly interested in how to align CS research with the SOI paradigm, rather than exploring alignment from the reverse perspective. To gain a comprehensive understanding of the current engagements of SOI scholars with fair AI, along with identifying constraints and facilitators, and possible alignment with CS, future research is needed.
@@
The third limitation pertains to the interview setting employed in our study. While qualitative interviews are valuable for exploring narratives, recounting stories, and uncovering constraints and challenges – aligning with our research interest – they also capture retrospective rationalizations of practices and events. Recognizing this, we encourage future research endeavors to conduct (more) ethnographic studies within the laboratory. Engaging in ethnographic studies allows investigating the doability of tasks in ‚real-time‘, observing activities as they unfold. We think that this will allow for a deeper understanding of the ambiguities of doability, and an exploration of the more tacit aspects of fair ML, which for those we interviewed might have already been ordinary and thus less noteworthy.
@@
Fourth, the advantage of examining researchers’ accounts in qualitative settings lies in providing interview participants with the opportunity to articulate ambivalence, correct their perspectives, and even pose questions in return. However, to expand generalizability, future research could broaden the scope of inquiry by incorporating quantitative surveys that could facilitate the collection of data from a larger, and ideally, a more diverse group of participants. Notwithstanding, future research could also consider introducing alternative ways to collect qualitative data, such as the elaboration of hypothetical scenarios (e.g., vignettes, speculative writing) in tandem with our interview partners to reflect upon different real-world situations. Fifth, we delineated two paradigms, but it is important to consider the possibility of additional paradigms existing, emerging and changing (Clarke and Star, 2008). This insight also extends to the actors within the social world and their alignments with the CS lab. Notably, investigating actors excluded from the fair AI world, e.g. identifying who is left out and why, promises deeper insights into unequally distributed power among the social world actors.
@@
Finally, our research findings offer a critical perspective on Fujimura’s theory of articulation work, suggesting a need to revisit some of its premises, as our findings shine a light on an important gap: the theory focuses on alignment with the social world, but much less on alignment between social world actors with differing and ambiguous stances. Complicating this situation, and as our findings underscore, there is an asymmetrical distribution of power among social world actors. We argue for a deeper exploration of power dynamics within articulation work to elucidate how some exert greater influence in shaping (CS) research agendas and production work.
@@ 8. Outlook
To understand fair AI, it is essential to shift the perspective towards its actors, their everyday practices, and their challenges. By investigating the conditions and constraints that CS researchers experience, as well as their frustrations, needs, and desires, we gain important insights into how fair AI is currently made doable, and can derive novel inspirations about how else it could be. This points to a need for the consideration of how to embrace more interdisciplinary and social-oriented research into our work and not be confined by what is deemed doable. It also requires exploring avenues that challenge the organizational (and epistemic) status quo.
As CSCW and STS theories remind us, scientific research is, fundamentally, a form of work. Of course, engaging in fair AI could also extend beyond the formal boundaries of work. This suggestion is not intended to replace integrating SOI into computational practices, but rather serves as a reminder of the diverse possibilities for meaningful engagement with fairness and social justice across various aspects of our lives (D’Ignazio and Klein, 2020; Benjamin, 2022), extending beyond professional responsibilities.
Some of the issues we have identified, and the proposals for addressing them, mirror long-standing struggles and some constructive developments in academic research more generally. For that reason, this current moment of fair AI research could be seen as a springboard to motivate progress in academic practices and knowledge production beyond the CS paradigm.
@@ Acknowledgements.
This work has received funding from the European Union’s Horizon 2020 research and innovation programme under Marie Sklodowska-Curie Actions (grant agreement number 860630) for the project “NoBIAS - Artificial Intelligence without Bias”. This work reflects only the authors’ views, and the European Research Executive Agency (REA) is not responsible for any use that may be made of the information it contains.

### v2

@@ 5.3. Openness to the SOI paradigm and tinkering
We conclude our findings by pointing out interview participants’ micro-strategies and attempts to integrate some SOI perspectives into their everyday research practices.
@@
Some of our interview participants, notably two professors, R4 and R5, both situated in German universities along with R8, the professor from an Italian university, and R10, a professor from a UK university, expressed their wish for fair AI research to become more inclusive of SOI perspectives. Some of them acknowledged their own agency and responsibility in attaining this. In particular, R10 emphasized how he tried to negotiate and shape the conditions of his CS lab to align with a more inclusive research approach. This, as he explained, included actively taking responsibility for conducting more research with interdisciplinary actors, such as scholars from the social sciences. As a senior researcher, he had experienced that such collaborations did not emerge spontaneously or by themselves but required proactive efforts by researchers, such as to “negotiate and have a more active role in your occupation to be able to keep the things that you’re happy with.” (R10)
@@
Some interview participants built their own accounts of doing SOI work by gradually deviating from usually performed CS tasks. Drawing on Fujimura (1987) and Knorr (1979), we understand these micro-strategies as tinkering. We especially observed tinkering with SOI methods. For instance, R7, the professor at the Dutch University who, as mentioned in section 5.1, engaged his PhD students in conducting “short ethnographic studies”, as he coined them, at field sites of their collaborative partners. At the same time, the ethnographic study was further adapted and tinkered with, so that articulation work did not increase significantly. The professor conveyed that in his lab, ethnography - which is usually conducted for a longer period of time, with the researcher becoming very familiar with the field, following rigorous strategies of self-reflection, ‘thick description’, and a back-and-forth between theory and empirical material (Hammersley and Atkinson, 2019; Spradley, 2016; Geertz, 1973) - was only conducted for several days, and without formal training. In another example, R1 told us a story about how he aimed to conduct a qualitative study to test users’ perceptions of an algorithmic system he and his colleagues were designing. However, to avoid additional articulation work (e.g. setting up a sampling strategy, reaching out to interview participants, organizing the interview setting, etc.), he turned to his colleagues as proxies for the intended users of the algorithmic system.
@@
Our findings show that there was a desire to accomplish impactful (fairness) research that coexisted with varying degrees of integration of SOI work into everyday research practices. Even researchers with a more critical stance towards the SOI paradigm demonstrated some (surface-level) engagement with it. Yet, due to the organizational constraints of the CS lab, SOI work was adjusted or modified, e.g. tinkered with, to sidestep additional articulation work.
@@ 6. Discussion
Bringing Fujimura’s study in dialogue with our findings, we now discuss the gap between the CS and SOI paradigms as a problem of ‘organizational alignment’.
@@
Our first argument is that computational researchers focus on the CS paradigm of fairness largely “because it works”, as to quote Knorr-Cetina (1981). Conversely, adopting the SOI paradigm necessitates acquiring new material resources and staff, reconfiguring publication venues, and dedicating additional time to novel methodological processes.
@@
In our study, we identified similar resources as crucial for making the CS paradigm of fairness research work. Much like in Fujimura’s observations, computational researchers pragmatically leveraged resources for fairness research just as they would for other research problems. This included acquiring funding and utilizing technologies, staff, and skills in ways similar to how they approached “other ML problems” (R5). What partly set fairness research apart from other ML problems, though, was its particular reliance on societal data that presented compelling societal use cases (or ‘real-world scenarios’). Computational researchers encountered challenges in legitimizing fairness research without access to such data, highlighting that technical resources shaped what was considered a meaningful advancement in the fair AI field by acting as facilitators.
@@
As Fujimura (1987) highlights, uncertainty increases articulation work as it constrains researchers’ possibilities to plan. In our case study, uncertainties posed constraints on computational researchers. The ‘projectification’ of research, with limited time frames, created general challenges for researchers in planning for the long-term. Information overload (e.g. ‘noise’) added to these uncertainties, as researchers were challenged to decide on their own accord which contributions to the field were important enough to extend, incorporate, or reference in their work. In that regard, the lack of regulatory and technical standards (e.g. benchmarks, toolkits, metrics) for fairness also increased uncertainty. As a result, it required more articulation work to find individualized solutions with the collaborative partners of computational researchers.
@@
If uncertainties and burdensome articulation work prevented computational researchers from fully engaging with the SOI paradigm, how can SOI become a ‘doable problem’ for them? From our theoretical standpoint, we now examine this question for each of the three organizational levels.
@@
The doability of a research problem greatly relies on the positions of the social worlds. Fujimura (1987) and Clarke and Star (2008) have noted that at times, there may be multiple intersecting social worlds, each with its own interests. This can increase the complexity of the articulation work required. As previously discussed, the fair AI world faces new emerging actors, such as regulatory bodies, becoming more important. Furthermore, different interpretations within social worlds exist regarding what is considered doable. This does not necessarily imply that fairness loses credibility because it is “empty” (Laclau, 2001; MacKillop, 2018; Kalluri, 2020), but that it can rather hold varying significance for different individuals (e.g., a so-called ‘boundary object’ (Star and Griesemer, 1989)). From this perspective, the much-requested call for standards, as sometimes indicated by our respondents, could potentially hinder interdisciplinary collaborations and alignment with diverse social worlds. Perhaps instead of such ethical or regulatory standards on fairness for computational researchers to translate into local (lab) practices, we need more alignment between different social worlds (e.g. CS community, regulatory bodies, companies, marginalized communities). This could include explicit stances from the social worlds’ actors acknowledging the SOI paradigm of fairness and making the SOI perspective an integral part of what is deemed as “success” (Knorr-Cetina, 1981).
@@
On the level of the laboratory, the criteria of how research is evaluated, what counts as a successful publication, and which venues are treated as important, could be re-evaluated. The success of research is still mostly defined via publications (and within these, the ‘weight’ relies on rigid and decontextualized hierarchies of venue quality). The quest for other forms of visibility is, and will remain, an open problem for both scientometrics theory and academic institutions’ practices. Academic institutions’ ongoing developments of evaluation and indicator systems (which are, to some extent, inherently self-defeating) continues141414e.g., KU Leuven’s position on the responsible use of metrics in the assessment of research(ers).. One potential approach to follow is the purposeful inclusion of SOI-related goals into project proposals, so that project scope and timeline explicitly account for research initiatives such as thorough user studies, to give an example. Following that, it is also important to account for the fact that an enhanced SOI-focus for a project will likely also require a more prominent role for actors from the SOI paradigm within the project structure.
@@
We have made the point that much of what drives the current situation is a need to limit articulation work for overburdened (computational) researchers. Therefore, we acknowledge that suggesting changes on the individual or experimental level could actually increase articulation work. However, we need alternative strategies that individual researchers can take if they wish to make impactful changes to the kind of research that gets done. One strategy that we propose is taking a conscious awareness of how we allocate time to various tasks and a thoughtful consideration of how tasks are driven by optimizing productivity and efficiency. From our own experience in collaborating for this paper, we would like to emphasize that interdisciplinary alignment takes a considerable amount of time and that there are misses along the way. To navigate such phases successfully, requires, as one of our interview partners also pointed out, to allow our research to be guided by excitement and curiosity even if that initially seems undoable.
@@
Paradoxically, making SOI research doable could also accentuate, or create, new hierarchies between paradigms and disciplines. Resources are inherently unequal across disciplines and academic institutions (Merton, 2000; Larregue, 2018; Albert et al., 2009). As supported by our findings, this can lead to a scenario where SOI methods and theories are gradually absorbed by computational disciplines that are already situated in a more powerful and financially advantaged position, particularly when it comes to ML research projects. In configuring interdisciplinary projects, close attention to fair resource and influence distributions amongst disciplines and institutions is required. Additionally, there is a need to address open questions about how to make something doable without perpetuating existing hierarchies.
@@ 7. Limitations and Future Research
We hope that our study opens some novel avenues for investigating fair AI from the perspective of the CS researchers involved in it. Yet, our study is limited to only a small group of researchers. While we were especially interested in the perspectives and experiences of these computational researchers, this reduces the scope of our work and hence limits the generalizability of the claims made regarding CS laboratories at large. Future research could integrate the perspectives of other relevant actors in the laboratory. Beyond CS, the doability of fair AI depends on the collaborative efforts of engineers, data scientists, ethicists, research group leaders, or IT experts who may be all situated in the (same) laboratory. Further research could investigate these micro-level actors’ interactions with each other, as well as with meso-level actors more closely.
@@
Second, we would find it interesting to further systematically explore our interviews regarding our respondents’ definitions of “fair AI” compared to a rather more flexible and open definition adopted while talking about their ongoing research projects and professional routines (Kinder-Kurlanda and Fahimi, [n. d.]). Related to that, we were explicitly interested in how to align CS research with the SOI paradigm, rather than exploring alignment from the reverse perspective. To gain a comprehensive understanding of the current engagements of SOI scholars with fair AI, along with identifying constraints and facilitators, and possible alignment with CS, future research is needed.
@@
The third limitation pertains to the interview setting employed in our study. While qualitative interviews are valuable for exploring narratives, recounting stories, and uncovering constraints and challenges –- aligning with our research interest –- they also capture retrospective rationalizations of practices and events. Recognizing this, we encourage future research endeavors to conduct (more) ethnographic studies within the laboratory. Engaging in ethnographic studies allows for investigating the doability of tasks in ‘real-time’, and observing activities as they unfold. We think that this will allow for a deeper understanding of the ambiguities of doability, and an exploration of the more tacit aspects of fair ML, which for those we interviewed might have already been ordinary and thus less noteworthy.
@@
Fourth, the advantage of examining researchers’ accounts in qualitative settings lies in providing interview participants with the opportunity to articulate ambivalence, correct their perspectives, and even pose questions in return. However, to expand generalizability, future research could broaden the scope of inquiry by incorporating quantitative surveys that could facilitate the collection of data from a larger, and ideally, a more diverse group of participants. Notwithstanding, future research could also consider introducing alternative ways to collect qualitative data, such as the elaboration of hypothetical scenarios (e.g., vignettes, speculative writing) in tandem with our interview partners to reflect upon different real-world situations.
Fifth, we delineated two paradigms, but it is important to consider the possibility of additional paradigms existing, emerging, and changing (Clarke and Star, 2008). This insight also extends to the actors within the social world and their alignments with the CS lab. Notably, investigating actors excluded from the fair AI world, e.g. identifying who is left out and why, promises deeper insights into unequally distributed power among the social world actors.
@@
Finally, our research findings offer a critical perspective on Fujimura’s theory of articulation work, suggesting a need to revisit some of its premises, as our findings shine a light on an important gap: the theory focuses on alignment with the social world, but much less on alignment between social world actors with differing and ambiguous stances. Complicating this situation, and as our findings underscore, there is an asymmetrical distribution of power among social world actors. We argue for a deeper exploration of power dynamics within articulation work to elucidate how some exert greater influence in shaping (CS) research agendas and production work than others.
@@ 8. Outlook
To understand fair AI research, it is essential to shift our perspective toward its actors, their everyday practices, and their challenges. By investigating the conditions and constraints that CS researchers experience, as well as their frustrations, needs, and desires, we gain important insights into how fair AI is currently made doable, and can derive novel inspirations about how else it could be. This points to a need for the consideration of how to embrace more social-oriented and interdisciplinary research in our work and not be confined by what is deemed doable. It also requires exploring avenues that challenge the organizational (and epistemic) status quo.
As CSCW and STS theories remind us, scientific research is, fundamentally, a form of work. Some of the issues we have identified, along with the proposals to address them, mirror long-standing struggles and constructive developments within academic work more broadly. However, engaging in fair AI could also reach beyond the formal boundaries of work. This suggestion is not meant to replace integrating SOI into computational practices but rather serves as a reminder of the diverse opportunities for meaningful engagement with fairness and social justice across various aspects of our lives (D’Ignazio and Klein, 2020; Benjamin, 2022), surpassing professional responsibilities.
@@ Acknowledgements.
This work has received funding from the European Union’s Horizon 2020 research and innovation programme under Marie Sklodowska-Curie Actions (grant agreement number 860630) for the project “NoBIAS - Artificial Intelligence without Bias” and from the Flemish Government (AI Research Program). Maria-Esther Vidal is partially supported by Leibniz Association, program “Leibniz Best Minds: Programme for Women Professors”, project TrustKG-Transforming Data in Trustable Insights; Grant P99/2020. This work reflects only the authors’ views, and the European Research Executive Agency (REA) is not responsible for any use that may be made of the information it contains.